OpenStudy (anonymous):

A fleet of nine taxis is to be dispatched to three airports in such a way that three go to airport A, five go to airport B, and one goes to airport C a) If exactly one of the taxis is in need of repair, what is the probability that it is dispatched to airport C? I got 1/9 as my answer . Using multinomial coefficients b) If exactly three of the taxis are in need of repair, what is the probability that every airport receives one of the taxis requiring repairs? I got 15/504 . Are these answers right?

OpenStudy (anonymous):

do for a) i did \[\left(\begin{matrix}8 \\ 3\hspace{2mm}5\hspace{2mm} 0\end{matrix}\right) =\frac{8!}{3!5!0!}=56\] and then this over the total number of ways = 504 ( obtained the same way but 9! instead of 9